A 5.5 g glass bead is charged by the removal of 10 x 10^10 electrons. What electric field strength (in units of N/C) will cause the bead to hang suspended in the air?
Answer #1

Charge of one electron = 1.6e-19 C

Charge of 10e10 electrons =10* 1.6e-9 C

Force acting on the charge = EQ where E is the electric field

Gravitational pull = mg = 0.00055 kg* 9.8m/s^2 = 0.00539 N

EQ= mg

10*1.6e-9*E = 0.00539

E =4.96e+5 (N/C) or (V/m)


answered by: lotus